Calculate the number of molecules of bromine in a 2.97×105 g sample of bromine.
1
Expert's answer
2019-10-21T07:16:28-0400

The amount of bromine molecules in moles is:

"n(Br_2)=\frac{m(Br_2)}{M(Br_2)}=\frac{N(Br_2)}{N_A}"

So the number of molecules of bromine can be found:

"N(Br_2)=\frac{m(Br_2)}{M(Br_2)} \cdot N_A=\frac{2.97 \times 10^5g}{160\frac{g}{mol}} \cdot 6.02 \times 10^{23} mol^{-1}=1.12 \times10^{27}"